OpenStudy (exoexoexo):

simplify and express with positive exponents (xy)^4/(xy)^3.(0.2)^-3

OpenStudy (exoexoexo):

\[\left( xy \right)^{4} \div \left( xy \right)^{3} \times \left( 0.2 \right)^{-3}\]

OpenStudy (tkhunny):

Have you considered simply collecting? Start with \(x^{Stuff}\) What goes in the "Stuff" zone for x?

OpenStudy (exoexoexo):

I simplified the x and y into \[x ^{4}y ^{4}\] and \[x ^{3}y ^{3}\]

OpenStudy (tkhunny):

\(x^{4-3} = x^1 = x\) You do the y's.